Ecological Water Filitration Inc is going to issue a $400,000 face value of 10%, 15 year bonds. The bonds are dated June 30, 2017, call for semi annual interest payments and mature on June 30, 2032. a. Compute the price investors should offer if they seek a yield of 8% on these bonds. Also, computer the first 6 months interest, assuming the bods are issued at this price. Use the interest method and calculate all amounts to the nearest dollar. b. Repeat part a. assuming investors seeks a 12% yield.
Answer #1

Answer a:

Face value = $400,000

Semiannual interest payment = 400,000 * 10% / 2= $20,000

Number of semiannual periods = 15 * 2 =30

Semiannual yield required = 8%/ 2 = 4%

To get price, we will use PV function of excel:

= PV (rate, nper, pmt, fv, type)

= PV (4%, 30, -20000, -400000, 0)

= $469168.13

Price investor should offer = $469,168

Interest expense for the first 6 months = $469,168.13 * 8%/2 = $18,766.73

Interest expense for the first 6 months = $18,767

Answer b:

Assume investor seeks a 12% yield.

Price = PV (12%/2, 30, -20000, -400000, 0) = $344,940.67

Price investor should offer = $344,941

Interest expense for the first 6 months = $344,940.67* 12%/2 = $20,696.44

Interest expense for the first 6 months = $20,696